OpenStudy (anonymous):

Write the equation of the line that passes through each pair of points in slope-intercept form. (0, -6) and (3, 15)

OpenStudy (anonymous):

The gradient is difference in y coordinates / difference in x-coordinates so 15 - (-6) / 3 -0 = 21 / 3 = 7. Then if equation of a line is y = mx + c, then rearranging c = y - mx choose your any point for the y and x coordinates, for example (3,15) to find the y-intercept. so if m=7 x = 3 y = 15 and equation c = y-mx what is c?

OpenStudy (anonymous):



OpenStudy (anonymous):

Have you understood ?

OpenStudy (anonymous):

lol why write it all out again? :D